OpenStudy (anonymous):

Identify the y-intercept in this linear equation 3x+y=6 A. 2 B. 3 C. -3 . D. 6

OpenStudy (anonymous):

\(y\)-intercepts are always found by substituting in \(x=0\) and solving for \(y\).

OpenStudy (anonymous):

So is the answer 6?

OpenStudy (anonymous):

Yes, D so it would look like this if x=0 (0,6)

OpenStudy (anonymous):

3x+y =6 3(0) + y =6 y = 6 Yes.

OpenStudy (anonymous):

@Ashley1997 which inequality is equivalent to 4x-3(x-2)>24 A. 7x+6>24 B. 4x-3x+6>24 C. 4x-3x-2>24 D. X-6>24

OpenStudy (anonymous):

@thechocoluver445

OpenStudy (anonymous):